OpenStudy (anonymous):

Find K \[(2k-2)^2-4(-4k+36)=0\]

OpenStudy (anonymous):

It's OK. I solved it!!

OpenStudy (ash2326):

We have \[(2k-2)^2-4(-4k+36)=0\] We could expand and solve for k. But we'll try a different method Let's rewrite the equation \[(2k-2)^2-4*(-2)(2k-18)=0\] or [(2k-2)^2+8(2k-2-16)=0\] Let's substitute 2k-2=r \[r^2+8(r-16)=0\] now \[r^2+8r-16=0\] Let's add 16 and subtract 16 \[r^2+8r+16-32=0\] now \[(r+4)^2-32=0\] or \[r+4=\pm \sqrt {32}\] r=2k-2 \[2k+2=\pm 4\sqrt {2}\] we get \[k=-1\pm 2\sqrt{2}\]

OpenStudy (anonymous):

That's not the right answer, but thanks for trying anway. I have the answer :D

OpenStudy (ash2326):

What answer you have? I'll recheck

OpenStudy (anonymous):

k= -7 or 5

OpenStudy (ash2326):

Got my mistake. :P

OpenStudy (anonymous):

great :D

OpenStudy (ash2326):

\[r^2+8(r-16)=0\] \(r^2+8r-128=0\) (this is where I made mistake earlier) add and subtract 16 \[r^2+8r+16-144=0\] \[(r+4)^2=144\] \[r+4=\pm 12\] r=2k-2 \[2k-2=\pm 12\] \[2k=2\pm 12\] \[k=1 \pm 6\] \[k=-5 or 7\] :D

OpenStudy (anonymous):

Yeah, got it :D